FEMTOSECOND STUDIES OF HUMAN THIOREDOXIN: DISSECTION OF COMPLEX DYNAMICS IN PROTEIN

\underline{WEIHONG QIU}, DONGPING ZHONG, OSU Biophysics Program, The Ohio-State University,Columbus, Ohio-43210, USA; WENYUN LU, Department of Physics,The Ohio-State University, Columbus-43210,USA.

Protein dynamics are complex processes and separation of these intertwined dynamics has been a great challenge to experimentalists. By integrating femtosecond up-conversion technique and site-directed mutagenesis, we are able to dissect this complex dynamics into elementary processes. Here we report our first studies of the complex dynamics in human thioredoxin. We separated the change-transfer process between Trp and the disulfide-bond from the hydration dynamics on the protein surface, both of which were found to occur on the similar time scale of 20 picoseconds.